Pharmax Distributors specializes in the distribution of pharmaceutical products at the retail level. Pharmax's customer base is comprised of three main retail groups: general supermarket chains, large drugstore chains, and small, single-store "Ma & Pa" pharmacies. To evaluate the relative profitability of each customer group, Pharmax plans to use an allocation system that recognizes the relative demands that these groups place on the company's resources. Parmax's chief accountant examined the firm's support costs and activities and collected the following data for a typical operating period: Support Activity Budgeted Cost Total Cost Driver Order processing $75,000 1,910 customer orders Line item processing $65,000 19,620 line items Store deliveries $68,000 1,470 deliveries Carton handling $76,000 67,050 cartons shipped She also gathered the following data for a typical operating period: General Supermarket Drugstore Chains Ma & Pa Customer orders 130 380 1,400 Line items 1,820 3,800 14,000 Deliveries 100 270 1,100 Cartons shipped 27,000 20,250 19,800 REQUIRED [ROUND ALL ALLOCATION RATES TO THE NEAREST CENT AND ALLOCATIONS TO THE NEAREST DOLLAR.] Part A (5 tries; 5 points) If Pharmax uses a single-driver system to allocate its support costs, with number of customer orders as the single driver, how much would be allocated to its Ma & Pa customers? Tries 0/5 Part B (5 tries; 5 points) If Pharmax uses an activity-based costing system to allocate its overhead costs, how much would be allocated to its Drugstore Chain customers?
Answer #1

a)

Activity Total cost
Order processing 75000
Line item processing 65000
Store deliveries 68000
Carton handling 76000
Total cost 284000
Total activity (Customer order) 1910
Activity rate [284000/1910] $ 148.69 per customer order
Cost allocated to MA & PA customers (actual activity used *activity rate) 1400 *148.69 = $ 208166

b)

COST ALLOCATED TO DRUGSTORE CHAIN CUSTOMERS
Activity Total cost [A] Estimated activity [B] Activity rate [A/B] ACTUAL activity cost allocated (Actual activity *activity rate)
Order processing 75000 1910 75000/1910=$ 39.27 Per customer order 380 14922.6
Line item processing 65000 19620 65000/19620 = $ 3.31 per line item processed 3800 12578
Store deliveries 68000 1470 68000/1470 =$ 46.26 per deliveries 270 12490.2
Carton handling 76000 67050 76000/67050 =$ 1.13 per carton shipped 20250 22882.5
TOTAL COST ALLOCATED TO DRUGSTORE CHAIN CUSTOMERS 62873